OpenStudy (anonymous):

solve for x 2In =(x+9) 0.4=log x 1=In(x+7) you invest 4,100 in an account that pays an interest rate of 7.5% compounded continuously. Calculate balance after of your account after 8 years? You invest 1,200 in an account that pays an interest rate of 7.25% compounded continuously calculate balance after 7 years.

OpenStudy (anonymous):

instead of in put ln can you edit the question pls

OpenStudy (anonymous):

0.4=log x \[0.4=\log_{10}x \implies 10^{0.4}=x\]

OpenStudy (anonymous):

\[1=\ln(x+7) \implies e^1=e^{\ln(x+7)}\] use \[e^{\ln b}=b\]

OpenStudy (anonymous):

I did edit it but it stills shows upas in